# What You See Is Not Always What You Get: An Empirical Study of Code Comprehension by Large Language Models

### v2

@@
Abstract.
Recent studies have demonstrated outstanding capabilities of large language models (LLMs) in software engineering tasks, including code generation and comprehension. While LLMs have shown significant potential in assisting with coding, it is perceived that LLMs are vulnerable to adversarial attacks. In this paper, we investigate the vulnerability of LLMs to imperceptible attacks, where hidden character manipulation in source code misleads LLMs’ behaviour while remaining undetectable to human reviewers. We devise these attacks into four distinct categories and analyse their impacts on code analysis and comprehension tasks. These four types of imperceptible coding character attacks include coding reordering, invisible coding characters, code deletions, and code homoglyphs. To comprehensively benchmark the robustness of current LLMs solutions against the attacks, we present a systematic experimental evaluation on multiple state-of-the-art LLMs. Our experimental design introduces two key performance metrics, namely model confidence using log probabilities of response, and the response correctness. A set of controlled experiments are conducted using a large-scale perturbed and unperturbed code snippets as the primary prompt input. Our findings confirm the susceptibility of LLMs to imperceptible coding character attacks, while different LLMs present different negative correlations between perturbation magnitude and performance. These results highlight the urgent need for robust LLMs capable of manoeuvring behaviours under imperceptible adversarial conditions. We anticipate this work provides valuable insights for enhancing the security and trustworthiness of LLMs in software engineering applications.
Adversarial attack, Large language models, Code comprehension
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyeccs: Computing methodologies Artificial intelligenceccs: Software and its engineeringccs: Software and its engineering General programming languages
1. Introduction
The success of Large Language Models (LLMs) in natural language processing (NLP) has driven their widespread adoption across many different sectors, one of which is software development. The presented ability to understand and generate code snippets through LLMs makes them highly accessible, resulting in the integration into developers’ workflows, including code generation (Du et al., 2024), program repair (Zhang et al., 2023), and vulnerability detection (Zhou et al., 2024).
@@ 1. Introduction
Recently, there are several studies demonstrating the feasibility of text-based adversarial attacks against LLMs, targeting performance outcomes of NLP tasks (Boucher et al., 2022; Formento et al., 2021; Li et al., 2020). While these studies have successful featured the LLMs vulnerabilities, it remains unknown concerning the feasibility of these attacks in real-world scenarios.  (Dyrmishi et al., 2023) argues that for an adversarial attack to be practical against a text-based LLM, the generated text must be both valid and natural. For attacks that fail to maintain high semantic consistency or the adversarial examples appear artificially generated, the corresponding real-world effectiveness will be limited. Through a survey of 378 human participants, their study concludes that existing text-based adversarial attacks lack imperceptibility, making them impractical in real-world scenarios where human reviewers are involved.
To address this research gap, this paper presents novel imperceptible adversarial attacks on source code, specifically targeting LLMs performing code analysis and comprehension tasks. Our goal is to understand whether or not the LLMs’ behaviour will be altered given manipulated code snippet which is visually indistinguishable from its original version when viewed by a human. We define ‘true imperceptibility’ as a malicious crafted code snippet that is completely indistinguishable from its original form. Prior works, such as (Liu et al., 2021), introduce imperceptible attacks in NLP but fail to achieve true imperceptibility, as the manipulated prompts, while semantically and visually similar, are still detectable when placed side-by-side. Our work bridges the gap by evaluating the practicality and effectiveness of truly imperceptible character attacks in real-world coding tasks with LLMs.
Thus, in this work, our attack method introduces a group of perturbation methods in code snippets to maintain true imperceptibility. Specifically, we leverage the special Unicode specification characters that subtly affect visual rendering while preserving readability for human reviewers. Inspired by Boucher et al. (Boucher et al., 2022), we develop four distinct imperceptible attack strategies targeting code analysis and comprehension tasks, namely code reordering, invisible coding characters, code deletions, and code homoglyphs. We include different LLMs solutions for the attack performance evaluation. The efficacy of these attacks are assessed based on two key performance metrics in a large-scale code analysis task, which are model confidence and response correctness. Experiment results demonstrates a strong negative correlation between perturbation and performance outcomes, with some notable differences among these LLMs, which we will further elaborate in Section. 5.
@@
  • We propose a black-box imperceptible coding character attack framework that introduces four distinct attack strategies against the state-of-the-art LLMs solutions. The strategies includes code reordering, invisible coding characters, code deletions, and code homoglyphs, which effectively manipulate LLMs’ behavior while remaining visually imperceptible for code analysis and comprehension task.
@@
The remaining sections of the paper are as follows. Section 2 presents the motivation of this study, while Section 3 discusses the background and related works. We provide the details of the overall framework design and experiment details in Section 4, and the experimental results in Section 5. Section 6 discusses our findings with novel insights, and Section 7 summarizes the paper.
2. Motivation
LLMs have recently transformed different sectors with enhanced accessibility through natural language. However, with LLMs being increasingly integrated into critical applications, concerns about their security, robustness and vulnerability have grown. Adversarial attacks via prompt injection have been demonstrated to degrade LLMs’ performance. While existing works have investigated the vulnerability on LLMs, they primarily focus on typical NLP tasks, such as sentiment analysis (et al., 2009) or text classification (et al., 2022).
This work is motivated by the intersection of software testing and NLP, particularly for LLM assisted software development tasks. LLMs have demonstrated exceptional capabilities in tasks such as code generation (Chen et al., 2021) and code comprehension (Fang et al., 2023). However, their behaviour heavily rely on the prompts, which introduces novel vulnerabilities to attack vectors that exploit the discrepancy between human interpretation and LLMs’ comprehension. For example, there are different special characters that can manipulate text and symbol rendering from Unicode specification (Consortium, 2021), as depicted in Table 1. This may create a significant divergence between what is displayed to the user and what the computers ‘see’.
@@ 2. Motivation
Table 1. Imperceptible attack
As most users interact with LLM-integrated tools through the user interface (UI) rather than at a deeper technical level, ensuring robustness against attacks that manipulate UI-level representation is critical. Conducting a comprehensive investigation into these human-imperceptible attacks on large-language models has the aim of not only providing a better understanding of the internal mechanisms of LLMs in current use, but also in identifying mitigation methods for the development of future generations of LLMs.
Thus, we argue that what is displayed to a human can be significantly different to what the large language models ‘see’. These perturbations as introduced in the LLMs’ system present potential ways to alter the LLMs’ behaviour. To this end, it is crucial to assess whether and to what extent an LLM can still execute its assigned tasks correctly when imperceptible perturbations are leveraged. It helps in closing the gap between user interaction and system instructions. To this end, in this work, we systematically investigate this challenge by devising four categories of imperceptible coding character attacks against LLMs for code analysis task. We also explore the different performance presented in different LLMs.
3. Background
3.1. Large language models
In this section, we briefly review the transformer architecture firstly introduced in 2017 (Vaswani et al., 2017), which underpins most of the popular large language models including BERT, BART, and GPT. Before transformer, the state-of-the-art approach to sequence modelling and transduction problems was considered to be recurrent neural networks (RNN). However, since RNNs factor computation along symbol positions of the input and output sequences, they exhibit an inherently sequential nature, which prevents parallelization. Without parallelization, the length of the sequence lengths is limited by hardware memory constraints. Increasing computational efficiency can curtail this limitation, but the fundamental restriction of sequential computation remains.
The key innovation the Transformer model exhibits is relying entirely on an attention mechanism in lieu of a recurrent network. Simply put, an attention mechanism allows the model to focus on different parts of the input sequence when trying to determine which parts are the most relevant to the eventual output sequence. Unlike a recurrent network, attention mechanisms are highly parallelizable, meaning that Transformer-model LLMs scale well with large amounts of training data. Critically, the attention mechanism also allows the model’s predictions to be interpreted, through analysis of the weights placed on each part of the model.
3.2. Applications of LLMs
Potential applications of large language models are versatile and far reaching, with transformative potential across many different sectors.
In healthcare, Cascella et al. (Cascella et al., 2023) evaluated the feasibility of using ChatGPT in four areas: supporting clinical practice, scientific text production, misuse in medical research, and reasoning about public health topics. While ChatGPT showed impressive abilities in tasks related to categorization and text generation, its responses in comprehension and reasoning tasks showed stereotyping. Most importantly, the authors highlighted the ‘hallucination’ phenomenon, where ChatGPT will produce answers that appear believable but are factually incorrect or logically inconsistent.
Chen et al. (Chen et al., 2021) introduces Codex in 2021, a GPT language model fine-tuned on publicly available code found on GitHub. The paper studies the Python code-writing capabilities of the model, which include using docstrings as input to generate code. A docstring is simply a string that documents a specific segment of code, like a code comment (Goodger and van Rossum, 2001). Using docstrings as LLM input to write code is an intersection of a natural language processing task with a code generation task – a similar intersection to the intersection between natural language processing and software testing. The investigation into Codex revealed some limitations in regard to this functionality, including difficulty parsing docstrings that described long chains of operations. Special attention was drawn to the risk of misalignment between the model’s capabilities and the user’s intent.
A more recent work by Fang et al. (Fang et al., 2023) presents a systematic evaluation of the code analysis capabilities of several popular LLMs, including GPT-3.5 and GPT-4. Two datasets were constructed, one with clean code samples in popular programming languages, and another with obfuscated versions of the JavaScript code samples. The paper found strong evidence suggesting that obfuscation techniques impact the abilities of the LLMs in generating code explanations, as well as a poor ability to generate de-obfuscated code without special optimization. Inspired by (Fang et al., 2023), we further explore the topic of using LLMs for code analysis tasks, in which we pin dowm the relevant findings that obfuscation does negatively affect LLM performance outcomes, and that there are still many unexplored ways to optimize LLMs in the presence of obfuscated code.
3.3. Attacks against LLMs
Now that the applications of LLMs have been demonstrated with regards to programming tasks, this section will outline some existing text-based attack strategies.
Formeto et al. (Formento et al., 2021) introduced the novel attack technique named the Special Symbol Text Attack (SSTA) in 2021. This technique exploits the fact that special symbols like punctuation marks or dashes contain downstream task information that can affect the LLM in NLP tasks such as sentiment analysis. The key finding is in the demonstration that symbols are an attack vector for text-based prompt injection attacks. However, the adversarial examples generated from this technique are clearly visible as confounding artifacts to the human eye. This leads us to the next paper.
Li et al. (Li et al., 2020) proposed BERT-attack in 2020, a method to generate adversarial examples using BERT against itself. BERT is a transformer-based language model, notable for being a dramatic improvement over other contemporaneous models (Devlin et al., 2018). The BERT-attack method consists of two steps: identify the most important word in the input sequence which contributes to the target prediction and keep replacing them with semantically similar words until the target model is fooled. The study used human evaluation to confirm that the adversarial examples generated with this method maintained high semantic similarity and grammatical correctness when compared to the clean samples.
The final piece of the background research is Boucher et al. (Boucher et al., 2022) entitled ‘Bad Characters: Imperceptible NLP Attacks’ in 2022. This paper introduces a new class of human-imperceptible attacks against NLP systems that exploit the visual similarity – and sometimes sameness – of certain Unicode characters. Four classes of imperceptible attacks are defined: Reordering, invisible characters, deletions, and homoglyphs, and adversarial examples were generated for a variety of models. Findings demonstrated that this new class of attacks could significantly degrade model performance on task-appropriate benchmarks, could target attacks, and could slow down processing speed by at least a factor of two. The potential for attacks such as these is enormous, with the authors discussing real-world applications such as evading filtering mechanisms, inhibiting content moderation, and degrading search engine results. Most relevant to us, this class of attack generates adversarial examples that target and attempt to manipulate the misalignment between the user’s intent and the system’s instructions.
4. Framework and Experiment Design
Figure 1. Overall imperceptible coding character attack framework
@@ 4.1. Research questions
  • RQ1:
    To what extent do imperceptible perturbations within code correlate with LLMs’ performance outcomes in software testing scenarios?
  • RQ2:
    Does a correlation exist between imperceptible perturbations and LLMs’ ‘confidence’, and what implications does this have for software testing?
  • RQ3:
    How significantly do different imperceptible perturbation methods impact the responses of LLMs?
  • RQ4:
    How do advancements in recent LLMs development improve their robustness against imperceptible adversarial perturbations in software testing scenarios?
4.2. Methodology
Figure 1 is the overall attack framework. Firstly, we frame the research using a unified prompting strategy in this work, which consists of three key elements:
  1. (1)
@@ 4.2. Methodology
  2. (2)
    Description of code snippet
  3. (3)
    Question: ‘Does the provided code match the provided description? Answer with either Yes or No.’
Using the OpenAI Chat Completions API (cha, 2024), the LLM is prompted to answer that same question for different pairs of code snippet and description. Since this study focuses on evaluating the effect of code perturbations rather than the natural language perturbations, only the code snippets are perturbed, and the descriptions remain constant. The response is limited to a binary output, i.e. Yes/No, so that we can leverage ‘logprobs’ (log, 2024) for accurately assessing response confidence during data analysis. The entire response generation process is fully automated, requiring no human interaction. Additionally, we have particularly designed the experiments that eliminate the need for human review to determine the correctness of LLM-generated responses. In next section, we will provide the details of the attack strategies.
4.3. Attack taxonomy
@@
  • Code reordering: Character U+202E is used so the Unicode specification can support characters from languages that read from right-to-left (202, 2023). Inserting it at the head and tail of the reversed string means that to the human eye, the perturbed section is rendered in the correct order, but the input will be the reversed string.
@@
  • Code homoglyphs: Homoglyphs are unique character that render to the same glyph. The Unicode Consortium publishes a document with the Unicode Security Mechanisms technical report (int, 2023) which details a mapping of characters that are intended as homoglyphs within the Unicode specification. These characters are intended to be mapped to the same glyph in font implementations and are thus ‘true’ homoglyphs. The technical report also includes a document which details a set of characters that are commonly visually confused, although these glyphs are not intended to be mapped to the same glyph in font implementations (uni, 2023). Since this second set is dependent on the choice of font implementation, and the extent to which they are visually similar is subjective, they were not considered for this study.
4.4. Attack generation
To increase the control of the experiment, the decision was made to create perturbed samples programmatically, rather than generatively. This means that the data perturbation process was open-source and consistent, rather than generating them using a process that both involves closed-source software and produces inconsistent output.
The perturbation process could not be exactly replicated across each category of perturbation, but each process shared in common the following 5 traits:
  1. (1)

### v3

# What You See Is Not Always What You Get: Evaluating GPT’s Comprehension of Source Code

@@
Abstract
Recent studies have demonstrated outstanding capabilities of large language models (LLMs) in software engineering tasks, including code generation and comprehension. While LLMs have shown significant potential in assisting with coding, LLMs are vulnerable to adversarial attacks. In this paper, we investigate the vulnerability of LLMs to imperceptible attacks. This class of attacks manipulate source code at the character level, which renders the changes invisible to human reviewers yet effective in misleading LLMs’ behaviour. We devise these attacks into four distinct categories and analyse their impacts on code analysis and comprehension tasks. These four types of imperceptible character attacks include coding reordering, invisible coding characters, code deletions, and code homoglyphs. To assess the robustness of state-of-the-art LLMs, we present a systematic evaluation across multiple models using both perturbed and clean code snippets. Two evaluation metrics, model confidence using log probabilities of response and response correctness, are introduced. The results reveal that LLMs are susceptible to imperceptible coding perturbations, with varying degrees of degradation highlighted across different LLMs. Furthermore, we observe a consistent negative correlation between perturbation magnitude and model performance. These results highlight the urgent need for robust LLMs capable of manoeuvring behaviours under imperceptible adversarial conditions.
I Introduction
The success of Large Language Models (LLMs) in natural language processing (NLP) has driven their widespread adoption across many different sectors, one of which is software development. The presented ability to understand and generate code snippets through LLMs makes them highly accessible, resulting in the integration into developers’ workflows, including code generation [Du24], program repair [Zhang23], and vulnerability detection [Zhou24].
@@ I Introduction
Recently, there are several studies demonstrating the feasibility of text-based adversarial attacks against LLMs, targeting performance degradation of NLP tasks [Boucher22, Formento2021, Li20]. While these studies have successfully featured the LLMs vulnerabilities, it remains unknown concerning the feasibility of these attacks in real-world scenarios. Dyrmishi et al. argues that for an adversarial attack to be practical against a text-based LLM, the generated text must be both valid and natural to human reviewers [Dyrmishi23]. For attacks that fail to preserve high semantic consistency or produce text that appears artificially generated, the corresponding real-world effectiveness will be limited. Through a survey of 378 human participants, their study concludes that existing text-based adversarial attacks lack imperceptibility, limiting their effectiveness in real-world scenarios where human reviewers are involved.
To address this research gap, this paper presents a novel class of imperceptible adversarial attacks on source code, specifically targeting LLMs in code analysis and comprehension tasks. Our goal is to investigate how the LLMs’ behaviour changes when presented with subtly manipulated code that appears identical to human reviewers. We define “true imperceptibility” as adversarial code snippet that is visually and semantically indistinguishable from its original form. While prior work such as [Liu21] explores imperceptible attacks in NLP, those manipulations often leave detectable artifacts while may be semantically and visually similar. Our work demonstrates the feasibility and effectiveness of truly imperceptible, character-level attacks in realistic software engineering scenarios, bridging the gap between theoretical vulnerabilities analysis and practical threats.
Thus, in this work, our attack method introduces a set of perturbation techniques to source code with the goal of achieving true imperceptibility. Specifically, we leverage the special Unicode specification characters that subtly affect visual rendering while preserving readability for human reviewers. Inspired by Boucher et al. [Boucher22], we design four distinct imperceptible attack strategies targeting code analysis and comprehension tasks, namely code reordering, invisible coding characters, code deletions, and code homoglyphs. We include different LLMs solutions, specifically with the state-of-the-art GPT models, for the attack performance evaluation. The efficacy of these attacks are assessed based on two key performance metrics in a large-scale code analysis task, which are model confidence and response correctness. Experiment results demonstrates a strong negative correlation between perturbation and performance outcomes, with some notable differences among these LLMs, which we will further elaborate in Section IV.
@@
  2. 2.
    We propose a black-box imperceptible coding character attack framework that introduces four distinct attack strategies against the state-of-the-art LLMs solutions. The strategies includes code reordering, invisible coding characters, code deletions, and code homoglyphs, which effectively manipulate LLMs’ behaviour while remaining visually imperceptible for code analysis and comprehension task.
  3. 3.
@@
The remaining sections of the paper are as follows. Section II presents the motivation and background of this study. We provide the details of the overall framework design and experiment setup in Section III, and report the experimental results in Section IV. Section V discusses our findings with novel insights, and Section VI summarises the paper.
II Motivation and Background
II-A LLM for Coding Tasks
Large Language Models (LLMs) have achieved strong performance across a broad range of natural language processing tasks, including sentiment analysis [Ye09], text classification [Dogra22], and scientific writing. In high-stakes domains such as healthcare, LLMs have been evaluated for clinical reasoning, summarisation, and patient guidance [Cascella23]. Despite these advances, LLMs are prone to hallucination, producing fluent yet factually incorrect outputs, and are often deployed without formal guarantees of robustness or security.
The adoption of LLMs in software engineering is emerging, with tools like GitHub Copilot [^2] https://github.com/features/copilot and Amazon Q Developer [^3] https://aws.amazon.com/q/developer/ integrating LLMs for code generation and review. Codex, a GPT-based model trained on large-scale public code repositories, exemplifies this trend by generating Python code from docstrings and natural language input [Chen21]. Studies have also explored LLMs’ capacity to explain and analyse code snippets, which forms the basis for tasks like static code analysis and review automation [Fang24].
II-B Imperceptible Perturbations and Technical Gap
Although adversarial prompt attacks have been widely studied in natural language processing, including synonym substitution and syntactic rewrites [Ye09, Dogra22], these techniques are typically visible and less applicable to source code tasks where minor character-level changes can affect semantics. In software engineering, especially in code review or comprehension, perturbations that are invisible to humans but still affect machine interpretation pose a unique and underexplored threat.
Boucher et al. [Boucher22] introduced a new class of attacks, termed imperceptible NLP attacks, which exploit the visual similarity or invisibility of certain Unicode characters. Their taxonomy included four categories: homoglyph substitutions, reordering of characters, invisible characters, and deletions. These perturbations degrade model performance on various NLP tasks without altering visible output. The study also demonstrated real-world implications, including bypassing content filters and degrading search engines, by misaligning human and machine understanding of the same string.
Table I illustrates typical examples adapted from the imperceptible attack class. For instance, right-to-left override (“U+202E”) and repeated backspace characters (“U+8”) are invisible in most editors but significantly alter the semantics of the input.
TABLE I: Examples of imperceptible attacks adapted from [Boucher22]
@@ II-B Imperceptible Perturbations and Technical Gap
While Fang et al.[Fang24] examined LLM behaviour under code obfuscation techniques such as variable renaming or restructuring, their work did not address character-level visual deception. Our work builds on the foundations laid by Boucher et al., extending the scope from NLP tasks to source code understanding and vulnerability analysis in LLMs.
II-C Real-World Attack Scenarios and Overtrust
Non-Expert Users. An increasing number of users, including hobbyists, students, scientists, copy and reuse code from LLM-generated outputs, online repositories, or technical forums. Lacking formal software engineering expertise, these users increasingly rely on the same LLMs (e.g., ChatGPT, Copilot Chat) to validate these code snippets [arxiv2401, acmNonExpert2024, danielZeroCode2023, arxiv2211insecure]. This creates a critical risk: perturbed or adversarially crafted code may be incorrectly interpreted as valid, with users unaware of the underlying syntactic or security implications of what they are copying.
Overtrust and Automation Bias. Many studies report that users tend to overtrust LLM outputs, even when flawed or speculative, [cacmCopilot2025, Fang24, overtrust2025, arxivTrustAI2024]. For instance, in healthcare, parents have shown a preference for ChatGPT responses over advice from certified physicians [kuParents2024]. In software engineering, such overtrust can manifest as the uncritical adoption of flawed code analyses or faulty model outputs.
Organisational Integration and Supply Chain Risk. We have seen enterprises are increasingly integrating LLMs into CI/CD pipelines and code auditing workflows, such as Amazon Q Developer and Copilot. If imperceptible attacks evade detection by LLM-powered validation layers, these malicious codes can be introduced into the software supply chain with internal security checks undetected [storek2025xoxo, infosecAICode2024, paloaltoAICyber2023, ibmAIOps]. Without adversarial hardening [dyrmishi2023empirical], the implicit trust placed in LLMs within development operations can become a significant security liability.
III Framework and Experiment Design
Figure 1: Overall imperceptible coding character attack framework
@@ III-A Research questions
  • RQ2: Does a correlation exist between imperceptible perturbations and LLMs’ “confidence”, and what implications does this have for software testing?
III-B Methodology
Figure 1 presents the overall attack framework used in this study. We adopt a unified prompting strategy composed of three key elements:
  1. 1.
@@ III-B Methodology
  2. 2.
    Description of the code snippet
  3. 3.
    A binary question: “Does the provided code match the provided description? Answer with either Yes or No.”
The OpenAI Chat Completions API [chatcomp] is used to submit each prompt to the target LLM. For each input, the code snippet is perturbed according to a specific strategy while the accompanying description remains unchanged. This setup isolates the effect of code-only perturbations on the model’s understanding and judgment.
The model’s response is constrained to a single-token binary answer: either “Yes” or “No”. This design choice ensures that the evaluation remains fully quantifiable and reproducible, avoiding the ambiguity inherent in open-ended responses. By enforcing a binary format, we enable direct, consistent comparison of correctness across all test cases and facilitate precise use of the “logprobs” property [logprobs] to measure model confidence. Additionally, this format mitigates the need for human interpretation, which would be required to judge multi-token explanations and could introduce subjective bias and inconsistency—particularly problematic in large-scale or automated evaluations. Finally, the Yes/No structure also reflects practical real-world scenarios, where users often pose concise security-related questions to LLMs, such as “Is this code safe?”
The entire response generation process is fully automated, requiring no human involvement for response validation or grading. To ensure scalability and objectivity, we designed the experiment to produce correctness and confidence scores programmatically, without relying on model self-evaluation or human annotation. The next section describes the specific perturbation strategies used to challenge the model’s comprehension capabilities.
III-C Attack taxonomy
@@
  • Code reordering: Character U+202E is used so the Unicode specification can support characters from languages that read from right-to-left [202e]. Inserting it at the head and tail of the reversed string means that to the human eye, the perturbed section is rendered in the correct order, but the input will be the reversed string.
@@
  • Code homoglyphs: Homoglyphs are unique character that render to the same glyph. The Unicode Consortium publishes a document with the Unicode Security Mechanisms technical report [intentional] which details a mapping of characters that are intended as homoglyphs within the Unicode specification. These characters are intended to be mapped to the same glyph in font implementations and are thus “true” homoglyphs. The technical report also includes a document which details a set of characters that are commonly visually confused, although these glyphs are not intended to be mapped to the same glyph in font implementations [unintentional]. Since this second set is dependent on the choice of font implementation, and the extent to which they are visually similar is subjective, they were not considered for this study.
III-D Attack generation
To increase the control of the experiment, the decision was made to create perturbed samples programmatically, rather than generatively. This means that the data perturbation process is open-source and consistent, rather than generating them using a process that both involves closed-source software and produces inconsistent output. The perturbation process could not be exactly replicated across each category of perturbation, but each process shared in common the following 5 traits:
  1. 1.